### v1

Abstract
Large Vision-Language Models (LVLMs) have shown significant progress in well responding to visual-instructions from users. However, these instructions, encompassing images and text, are susceptible to both intentional and inadvertent attacks. Despite the critical importance of LVLMs’ robustness against such threats, current research in this area remains limited. To bridge this gap, we introduce AVIBench, a framework designed to analyze the robustness of LVLMs when facing various adversarial visual-instructions (AVIs), including four types of image-based AVIs, ten types of text-based AVIs, and nine types of content bias AVIs (such as gender, violence, cultural, and racial biases, among others). We generate 260K AVIs encompassing five categories of multimodal capabilities (nine tasks) and content bias. We then conduct a comprehensive evaluation involving 14 open-source LVLMs to assess their performance. AVIBench also serves as a convenient tool for practitioners to evaluate the robustness of LVLMs against AVIs. Our findings and extensive experimental results shed light on the vulnerabilities of LVLMs, and highlight that inherent biases exist even in advanced closed-source LVLMs like GeminiProVision and GPT-4V. This underscores the importance of enhancing the robustness, security, and fairness of LVLMs. The source code and benchmark will be made publicly available. * indicates corresponding author.
Keywords:
Large Vision-Language Model Adversarial Visual-Instructions Bias Evaluation
1 Introduction
Over the past year, Large Language Models (LLMs) have achieved significant milestones, consistently demonstrating exceptional performance across a diverse range of natural language processing tasks. This success has spurred the development of LLM-based applications, reshaping our daily lives. More recently, alongside advanced closed-source Large Vision-Language Models (LVLMs) like GeminiProVision [1] and GPT-4V(ision) [2], many open-source LVLMs have emerged, such as Otter [3], InternLM-XComposer [4], ShareGPT4V [5], and Moe-LLaVA [6]. These open-source models propose various architectures and training methods to enhance the capabilities of powerful LLMs like Vicuna [7] and LLaMA [8], enabling them to understand images and perform multimodal tasks such as visual question answering [9], multimodal conversation [10], and complex scene comprehension [3]. Considering that LVLMs form the foundation for next-generation AI applications, addressing concerns related to their robustness, security, and bias is of utmost importance.
Figure 1: The overview of AVIBench.
LVLMs employ two input modalities, text and images, both susceptible to adversarial perturbations. While pioneering studies [11, 12] have assessed LLMs’ robustness against text-based attacks, there is a lack of specific exploration targeting LVLMs. Recent investigations of image attacks have examined limited LVLMs’ resilience against white-box attacks [13, 14], backdoor attacks [15], query-based black-box attacks [16], and transfer-based black-box attacks [17]. However, transfer-based black-box attacks rely on surrogate models to execute the attacks, posing challenges in finding an LVLM-agnostic surrogate model applicable to all LVLMs. White-box attacks, backdoor attacks, and query-based black-box attacks, which depend on the output probability distributions of LVLMs, may be impractical for online-accessed models, particularly closed-source LVLMs. Moreover, these attack methods may be constrained by their specific task design, such as image captioning [18] or visual question answering [13], thus limiting the evaluation’s comprehensiveness.
Figure 2: Comparison of LVLMs’ robustness of adversarial visual-instructions. The robustness of adversarial visual-instructions about (a) image corruptions, (b) decision-based optimized black-box image attacks, (c) text attacks, and (d) content bias attacks.
Furthermore, as the applications employing LVLMs continue to emerge, paying greater attention to the risks stemming from LVLMs’ inherent biases becomes imperative. These biases, influenced by factors such as gender, race, the propagation of unsafe information, and cultural influences, may erode user trust and undermine the credibility of the applications. It is important to emphasize that revealing model biases transcends the realm of technical challenges; it is also a moral imperative that cannot be overlooked.
This paper introduces AVIBench, a comprehensive benchmark designed to evaluate LVLMs’ robustness in the face of adversarial visual-instructions (text-image pairs), as illustrated in Fig. 1. AVIBench encompasses a diverse set of adversarial visual-instructions (AVIs) that target text and images. Specifically, we adapt LVLM-agnostic and output probability distributions-agnostic black-box attack methods to target LVLMs, resulting in a total of 10 types of text-based AVIs and 4 types of image-based AVIs. In addition to these attacks, we also introduce 9 types of content bias AVIs, addressing issues related to gender, violence, culture, racial biases, and more, to evaluate the biases inherent in LVLMs comprehensively. AVIBench is mainly constructed from Tiny LVLM-eHub [19], which is a benchmark for five categories of multimodal capabilities (nine tasks). Finally, we construct 260K AVIs for AVIBench, which, along with our open-source code, can be utilized as a convenient tool to evaluate LVLMs’ defense against AVIs. We evaluate a total of 14 different open-source LVLMs using AVIBench and present the results in Fig. 2. Additionally, we evaluate closed-source LVLMs, including advanced systems like GeminiProVision and GPT-4V, using content bias AVIs.
Through extensive experimentation and analysis of the evaluation results, we make several noteworthy findings (detailed in Section 4). This paper serves a dual purpose by establishing a significant benchmark for assessing the robustness of LVLMs and potentially inspiring the development of mitigation and defense methodologies within the research community. Our main contributions can be summarized as follows:
  • We introduce AVIBench, a pioneering framework and versatile tool for evaluating the robustness of LVLMs on AVIs. AVIBench is designed to accommodate various tasks, models, and scenarios.
  • AVIBench generates a comprehensive dataset of 260K AVIs spanning five multimodal capabilities and content biases. This extensive dataset serves as a stringent evaluation benchmark, systematically probing LVLMs’ defense mechanisms against AVIs.
  • We evaluate the abilities of 14 open-source LVLMs to resist adversarial AVIs and show extensive experimental results and findings, which also offer convenience for developing robust LVLMs.
  • We show that even advanced closed-source LVLMs like GeminiProVision and GPT-4V exhibit significant content biases. This finding underscores the importance of advancing research on secure and fair LVLMs.
2 Related Work
Large Vision-Language Models. LVLMs like Otter [3], InstructBLIP [20], PandaGPT [21], InternLM-XComposer [4], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6] have made significant progress in multimodal tasks. These models align visual features with textual information by leveraging knowledge from LLMs like Vicuna [7] and LLaMA [8]. Techniques such as cross-attention layers [23], Q-Former [24], one project layer [25], and LoRA [26] have been used to bridge the gap between language and vision. Given the crucial role of LVLMs in future user interfaces and multimedia systems, ensuring their robustness, security, and fairness is paramount. Our paper focuses on evaluating LVLMs’ resilience against AVIs.
Evaluation of Large Vision-Language Models. Recent advancements in LVLMs have led to improvements in datasets and evaluation methods. LVLM-eHub [27] and Tiny LVLM-eHub [19] organize multiple vision-language benchmarks, while MME Bench [28] introduces a new evaluation dataset. Other benchmarks like LAMM [29], MMBench [30], Seed Bench [31] also contribute to LVLM evaluation. However, there is a lack of research on assessing LVLMs’ resistance to attacks on text and image modalities. To address this, we propose a benchmark that evaluates LVLMs’ ability to withstand attacks on images, text, and bias.
Attacks for Large Vision-Language Models. Pioneering research explored various LVLM-specific image attacks on limited LVLMs, including white-box attacks [13, 14], backdoor attacks [15], query-based black-box attacks [16], and transfer-based black-box attacks [17]. However, white-box attacks, backdoor attacks, and query-based black-box attacks require knowledge of the model’s output probability distribution; transfer-based black-box attacks require finding a surrogate model that is difficult to obtain for all LVLMs. Thus, we are the first to adapt LVLM-agnostic and output probability distribution-agnostic decision-based optimized image attacks specifically tailored for LVLMs. We also incorporate image corruption as an attack method, covering 5 multimodal capabilities across 9 subtasks, while Zhang et al. [32] only utilize image corruptions to evaluate limited LVLMs on the image captioning task. To target the text inputs of LVLMs, we draw inspiration from black-box text attacks originally designed for LLMs[11], adapting and expanding them for LVLMs. Furthermore, although previous studies have explored gender bias in the outputs of LVLMs [33, 34], we reveal the inherent biases that exist in LVLMs by building more comprehensive content bias AVIs.
3 AVIBench
In this section, we first introduce the definition of Adversarial Visual-Instructions (AVIs), followed by an introduction to the key components of AVIBench: models, dataset, and the construction of AVIs.
3.1 Definition of Adversarial Visual-Instructions
Unlike adversarial examples that cause "misclassification," Adversarial Visual-Instructions (AVIs) contain intentionally designed images and text that specifically manipulate LVLMs’ behavior in a broader sense. AVIs are intentionally crafted by adversaries to induce incorrect, unsafe, and harmful behavior in LVLMs, aligning with the broader definition of adversarial examples in [35][11].
3.2 Models
We collect a total of 16 LVLMs, including 14 open-source models and 2 closed-source models, to create a model hub for evaluation. The open-source models consist of BLIP2 [24], LLaVA [36], MiniGPT-4 [25], mPLUG-owl [37], LLaMA-Adapter V2 [9], VPGTrans [38], Otter [3], InstructBLIP [20], PandaGPT [21], OpenFlamingo-V2 [23], InternLM-XComposer [4], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6]. The closed-source models are GeminiProVision [1] and GPT-4V(ision) [2]. To ensure a fair comparison, we carefully select the open-source LVLM versions with closely aligned parameters level. Details of these models can be found in the supplementary material.
3.3 Dataset
Our base dataset is derived from Tiny LVLM-eHub [19], which consists of 2,550 instructions and corresponding answers, and is organized into nine tasks that evaluate LVLM’s five multimodal capabilities: (1) Visual perception involves interpreting and understanding visual information. (2) Visual knowledge acquisition refers to the capability of LVLMs to acquire and understand visual information from images. (3) Visual reasoning involves the ability to reason and answer questions. (4) Visual commonsense measures the model’s comprehension of shared human knowledge about visual concepts. (5) Object hallucination refers to the phenomenon where LVLMs generate content that does not match the actual objects present in a given image.
Our AVIBench dataset generates 260K AVIs based on the base dataset. In particular, AVIBench includes 145,350 AVIs for image corruption, about 26,736 AVIs for decision-based optimized image attacks, 55,000 AVIs for content bias attacks, and 33,000 AVIs for black-box text attacks. For more comprehensive information, including details about the base dataset, evaluation criteria, and the specific composition of AVIs, please refer to the supplementary material.
3.4 Construction of Adversarial Visual-Instructions
This paper adapts the LVLM-agnostic and output probability distribution-agnostic black-box attacks to construct AVIs. The reason for utilizing these attack methods is that they solely rely on the LVLMs’ text response.
We denote the LVLM as $f_{\theta}$ and the dataset with $M$ visual instructions(i.e., image-text prompt pairs) as $\mathcal{D}=\{(\mathcal{I}_{m},\mathcal{P}_{m})\}_{m=1,..,M}$. The objective function of our AVIs’ construction is:
$\mathop{\arg}_{\delta_{\mathcal{I}},\delta_{\mathcal{P}}}\Gamma_{\{(\mathcal{I}_{m},\mathcal{P}_{m});\mathcal{G}_{m}\}\in\mathcal{D}}{Score}[f_{\theta}(\{(\mathcal{I}_{m}+\delta_{\mathcal{I}},\mathcal{P}_{m}+\delta_{\mathcal{P}});\mathcal{G}_{m}\})],$ (1)
where $\delta_{\mathcal{I}}$ and $\delta_{\mathcal{P}}$ represent the image perturbation and the text perturbation. $\mathcal{G}_{m}$ represents the ground truth annotations for the instruction $(\mathcal{I}_{m},\mathcal{P}_{m})$. Score denotes the model’s predicted score, which depends on the evaluation criteria for assessing the model’s multimodal capabilities (see our supplementary material). The functions $\mathop{\arg}$ and $\Gamma$ have distinct interpretations in different kinds of AVIs, which will be further explained in the subsequent sections.
3.4.1 Four Types of Image-based AVIs.
In Section 2, we explained the reasons for not using white-box attacks, backdoor attacks, query-based black-box image attacks, and transfer-based black-box image attacks. Consequently, the remaining image attack methods are limited to image corruption and decision-based optimized image attacks. These attacks focus on individual visual instructions, making $\Gamma$ negligible.
Image corruptions encompass a range of applied distortions, including noise, blur, weather effects, and digital distortions, to the images. It is vital to assess LVLM performance under these different corruption categories. In line with the methodology of Hendrycks et al. [39], we generate a corruption variant of our base dataset, comprising 19 corruption categories graded across three severity levels  [^1] We use three of the five levels in [39], namely 1, 3, and 5.. $\mathop{\arg}$ has no practical meaning here.
Decision-based optimized image attacks are widely utilized in image classification. We are the first to involve an adaptive modification of three state-of-the-art (SOTA) existing methods: PAR [40], Boundary [41], and SurFree [42] to attack LVLMs. Given the diverse subtasks and evaluation metrics for LVLMs’ natural language responses, we replace the original attack objective for image classification with "the score of the evaluation criteria for different tasks is 0." Specifically, an evaluation metric value of 0 indicates a successful attack. For instance, if the F1 score is 0, it signifies a successful attack on the Key Information Extraction (KIE) task. In this context, the symbol $\mathop{\arg}$ denotes $\mathop{\arg}min$.
3.4.2 Ten Types of Text-based AVIs.
To assess the robustness of LVLMs against text-based AVIs, we adapt the seven text attack methods mentioned in PromptBench [11]. They are organized by four levels of attacks, including character-level, word-level, sentence-level, and semantic-level attacks. Then, we further adapt three additional attacks: Pruthi [43], Pwws [44], Input-reduction [45] for more comprehensive evaluation. And $\mathop{\arg}$ also has no practical meaning here.
Character-level attacks contains TextBugger [46], DeepWordBug [47], and Pruthi [43]. Word-level attacks contains BertAttack [48], TextFooler [49], and Pwws [44]. Sentence-level attacks includes StressTest [50], CheckList [51], and Input-reduction [45]. Semantic-level attacks we employ are the same as that used in Zhu et al. [11].
In several subtasks, certain tasks feature distinct texts for each text-image pair. Our focus excludes these tasks from attack, directing our evaluation toward tasks where a common text segment exists in the instructions. For instance, in a Visual Question Answering (VQA) task with the instruction: "Question: Does this picture have symmetry? Choose the best answer from the following choices: yes no", the shared portion across all instructions is "Choose the best answer from the following choices:". This design aims to enhance the efficiency in demonstrating the efficacy of text-based AVIs.
$\Gamma$ represents the cumulative impact on all attacked instructions within the subtask of each multimodal ability. All attacks are adaptively modified based on the definition of scores across the different multimodal abilities. We employ ChatGPT to enhance the prompt and generate the three most effective prompts for each task to facilitate the attack, the detailed process is presented in supplementary material.
3.4.3 Nine Types of Content Bias AVIs.
Our analysis focuses on unsafe information, cultural bias, racial bias, and gender bias present within the LVLMs. We aim to gain insights into the model’s susceptibility to content bias attacks. And in content bias attacks, $\mathop{\arg}$ has no specific meaning. We ignore the original visual instructions $(\mathcal{I},\mathcal{T})$ and construct perturbed samples $(\delta_{\mathcal{I}},\delta_{\mathcal{T}})$ directly, which means we attack a combination of images and text and create a new dataset apart from the basic dataset, which consists of 55,000 instructions.
All attacked samples adopt the format of VQA. To generate these samples, we gather images of different induced types from the Internet and design corresponding questions. We then assess whether the LVLM exhibits prejudice by examining the model output in the form of a question.
Unsafe content bias attacks: we directly assess the LVLM’s capability to detect six categories of unsafe information: Harassment, Hate, Self-Harm, Sexual, Shocking, and Violence.
Culture content bias attacks: we explore weddings and funerals in diverse cultural contexts, encompassing China, Japan, Africa, and the United States. This aims to assess the LVLM’s accuracy in identifying these occasions.
Racial content bias attacks: we evaluate the LVLM’s performance across five dimensions: poverty, crime rate, wisdom, creativity, and art level for different races: Black, White, and Yellow.
Gender content bias attacks: we evaluate the appropriateness of male and female representations in ten different occupations: chef, dentist, doctor, lawyer, painter, police officer, scientist, singer, software developer, and teacher.
The specific process of image collection and visual instruction construction is detailed in the supplementary material.
Figure 3: Results of AVIs on the LLaVA-1.5. (a) Image corruption example. (b) Decision-based optimized black-box image attack example. (c) Black-box text attack example. (d) Content bias attack example.
4 Experiments
We present a visual demonstration of the robustness of 14$+$ LVLMs against various AVIs in Fig.2222The numerical values for the radar charts in Fig.1 can be found in the supplementary material.. Furthermore, Fig.3 displays some AVIs, offering a clearer understanding of our AVIs. For GeminiProVisionn [1] and GPT-4V [2], we only evaluated the robustness of them to content bias AVIs, as the access to these closed-source LVLMs is limited and restricted.
4.1 Evaluation Metrics
The Score in Equation 1 for individual subtasks are aligned with  [27, 19]. For image corruption in image-based AVIs and text-based AVIs, we use the Average Score Drop Rate (ASDR) as the evaluation metric, defined as follows:
$ASDR$ $=\frac{1}{M}\sum_{\{(\mathcal{I}_{m},\mathcal{P}_{m});\mathcal{G}_{m}\}\in\mathcal{D}}\frac{Score_{b}-Score_{a}}{{Score}_{b}},$ (2)
where $Score_{b}$ and $Score_{a}$ represent the $Score[f\theta({(\mathcal{I}_{m},\mathcal{P}_{m});\mathcal{G}{m}})]$ before and after attacks, respectively.
For decision-based black-box image attacks in image-based AVIs, we employ two evaluation methods: Attack Success Rate (ASR) and Average Euclidean Distance (AED) similar to [40]. ASR indicates the proportion of successful attacks. As the three black-box attacks use the same initial attack method [40], the ASR for all three attacks is identical. AED represents the average Euclidean distance between the image after a successful attack and the original image [^3] A smaller AED indicates a lower likelihood of the attack being detected by human eyes..
Regarding content bias AVIs, due to the design of each attack bias content, we evaluate the accuracy of VQA [27] for all content bias visual-instructions.
4.2 Results on Image Corruptions in Image-based AVIs
Table 1: Comparing the effectiveness of image corruptions in image-based AVIs. C-name, Gau., and Imp. refer to corruption name, Gaussian, and Impulse respectively. The best results are bold, the worst results are underlined.
C-name Fog Brightness Contrast Defocus Blur Elastic Frost Gau._Blur
ASDR 0.02 0.02 0.07 0.16 0.18 0.00 0.12
C-name Gau. Noise Glass Blur Imp._Noise JPEG Motion Blur Pixelate Saturate
ASDR 0.15 0.17 0.13 0.09 0.13 0.13 0.01
C-name Shot Noise Snow Spatter Speckle Noise Zoom Blur
ASDR 0.16 0.11 0.07 0.13 0.15
Table 2: Evaluation results of LVLMs’ robustness to decision-based optimized image attacks in image-based AVIs. R Ave. represents the average of the rows. Ave. ASR ($\downarrow$) and Ave. AED are calculated across five multimodal capabilities. Per., Kno., Rea., Com., and Hal. represent Visual Perception, Visual Knowledge Acquisition, Visual Reasoning, Visual Commonsense, and Object Hallucination respectively. P, B, and S correspond to PAR [40], Boundary [41], and SurFree [42] respectively. "-" indicates a task score of 0 before the attack.
BLIP2 In-BL. LA-V2 LLaVA MGPT m-owl Otter PGPT VPGT OF-2 In-XC. L-1.5 SGPT Moe R Ave.
Per. ASR 0.58 0.53 0.73 0.60 1.00 1.00 0.79 0.65 0.65 0.21 0.44 0.40 0.34 0.37 0.57
P 25.16 24.75 24.00 30.05 1.69 6.64 15.30 17.11 21.84 12.12 44.55 36.65 44.50 31.97 24.64
P+B 12.92 11.38 10.63 11.91 1.55 4.73 7.86 12.48 10.52 6.66 15.78 14.54 12.74 11.59 10.89
P+S 1.57 2.79 2.35 4.30 0.00 1.29 1.21 2.32 0.63 1.08 2.17 2.39 2.59 3.33 2.13
Kno. ASR 0.58 0.54 0.70 0.85 0.93 0.13 0.82 0.63 0.59 0.10 0.09 0.10 0.10 0.10 0.43
P 12.87 15.05 17.60 27.24 3.56 6.27 12.36 12.09 18.11 8.86 20.45 17.66 14.44 22.56 14.70
P+B 6.79 6.82 9.23 11.80 3.37 3.55 8.64 4.66 10.64 4.11 12.76 10.35 8.73 11.72 8.10
P+S 0.10 0.50 0.73 0.13 2.45 1.11 0.59 0.31 0.07 0.09 0.60 0.19 0.30 0.93 0.57
Rea. ASR 0.51 0.66 0.55 0.62 0.99 0.98 0.82 0.51 0.65 0.23 0.35 0.44 0.45 0.37 0.57
P 22.58 23.17 26.75 29.19 3.49 11.15 18.60 19.27 21.85 23.59 25.14 29.52 22.95 26.09 21.80
P+B 13.40 12.14 12.40 8.94 2.05 4.54 8.43 12.38 9.84 11.99 11.28 12.93 12.33 11.38 10.44
P+S 2.30 1.68 2.38 1.87 0.38 2.12 3.19 4.17 1.70 5.38 2.68 3.33 2.83 2.14 2.55
Com. ASR 0.33 0.37 0.24 0.45 0.65 0.94 0.56 0.25 0.36 0.17 0.38 0.33 0.36 0.29 0.40
P 38.99 26.03 26.54 31.93 8.74 7.65 24.51 27.17 24.75 22.13 26.21 34.58 47.88 36.57 27.43
P+B 15.64 14.76 9.86 6.07 7.18 4.67 11.68 15.00 13.30 10.98 14.13 8.36 11.19 14.46 11.51
P+S 8.79 7.61 7.56 4.15 2.44 3.12 6.11 9.11 8.92 4.81 7.76 4.23 4.68 5.26 5.94
Hal. ASR 0.37 0.44 0.43 - 1.00 0.99 0.33 0.11 0.66 0.01 0.68 0.83 0.80 - 0.55
P 30.50 32.71 37.70 - 0.21 33.02 36.42 81.90 30.26 0.18 38.49 53.33 45.92 - 35.05
P+B 17.83 10.99 12.36 - 0.03 6.82 9.99 11.27 16.42 0.00 16.28 16.12 17.94 - 11.34
P+S 9.17 5.77 3.99 - 0.00 4.73 5.11 7.26 0.04 0.00 8.10 8.49 11.44 - 5.34
Ave. ASR 0.47 0.51 0.53 0.51 0.91 0.81 0.66 0.43 0.58 0.14 0.39 0.42 0.41 0.28 0.50
Ave. AED 4.39 3.67 3.40 2.61 1.05 2.47 3.24 4.63 2.27 2.27 4.26 3.73 4.37 2.92 3.20
We assess the robustness of LVLMs against various image corruptions and present the experimental results in Fig. 1 (a). We observe that MiniGPT-4 [25] exhibits the strongest anti-corruption capability among the LVLMs, followed by Otter [3] and BLIP2 [24]. On the other hand, mPLUG-owl [37] shows the weakest performance, with an average performance drop of 17% across all image corruption attacks. Overall, all LVLMs are with average ASDR values consistently below 20%, which may be attributed to the availability of large-scale training data.
Table 1 compares different attack methods on all 14 open-source LVLMs. We find that Elastic, Glass_Blur, and Shot_Noise are more effective, with average ASDRs of 18%, 17%, and 16% respectively. On the other hand, Frost, Saturate, Fog, and Brightness are less effective, with average ASDRs of 0%, 1%, 2%, and 2% respectively. These findings can provide guidance to LVLM developers in designing targeted defense strategies.
4.3 Results on Decision-based Optimized Image Attack in Image-based AVIs
Table 2  [^4] In-BL., LA-V2, MGPT, m-owl, PGPT, VPGT, OF-2, In-XC., L-1.5, SGPT, Moe represent InstructBLIP, LLaMA-Adapter V2, MiniGPT-4, mPLUG-owl, PandaGPT, VPGTrans, OpenFlamingo-V2, InternLM-XComposer, LLaVA-1.5, ShareGPT4V, Moe-LLaVA, respectively. presents results for decision-based optimized image attacks in image-based AVIs. Regarding visual perception capability, MiniGPT-4 [25] and mPLUG-owl [37] are the most vulnerable LVLMs, while OpenFlamingo-V2 [23] exhibits high robustness (ASR: 21%). For visual knowledge acquisition capability, MiniGPT-4 [25] is the most vulnerable with an ASR of 93%, while InternLM-XComposer [4] performs well with an ASR of 9%. Other LVLMs, including mPLUG-owl [37], OpenFlamingo-V2 [23], LLaVA-1.5 [22], ShareGPT4V [5], and Moe-LLaVA [6], have ASRs below 15%, indicating satisfactory performance. In terms of visual reasoning capability, MiniGPT-4 [25] has the highest ASR of 99%, while OpenFlamingo-V2 [23] performs significantly better with an ASR 76% lower than MiniGPT-4. For visual commonsense capability, except for mPLUG-owl [37] (94%), the ASRs for other LVLMs are below 70%, with OpenFlamingo-V2 [23] being the best-performing LVLM. In evaluating object hallucination, OpenFlamingo-V2 [23] remains the top-performing LVLM, while MiniGPT-4 [25] exhibits the poorest performance, achieving a 100% ASR.
To summarize, MiniGPT-4 achieves the highest average ASR at 91%. Other LVLMs with notable performance include OpenFlamingo-V2 [23] at 14%, Moe-LLaVA [6] at 28%, and InternLM-XComposer [4] at 39%. Across various multi-modal capabilities evaluations, the three attack methods consistently demonstrate their effectiveness in the AED evaluation. Combining the PAR attack [40] with the Boundary [41] and Surfree [42] algorithms proves successful in reducing noise amplitude and improving attack detectability. Among the five abilities, visual perception and visual reasoning are the most vulnerable to attacks, with an average ASR of 57%, while visual common sense exhibits the least vulnerability, with an average ASR of 40%. These results emphasize the fragility of LVLMs and serve as a motivation for researchers to develop more robust LVLMs through targeted training approaches. It also highlights the need to enrich defense mechanisms to enhance their resilience against LVLM-agnostic and output probability distribution-agnostic black-box image attacks.
Table 3: Evaluation results of LVLMs’ robustness to text-based AVIs. Cha., Wor., Sen., Sem., Dee.ug, Inp.on represent character-level, word-level, sentence-level, and semantic-level, DeepWordBug, Input-reduction respectively. ASDR ($\downarrow$) is the metric. The best attack category is in bold, the worst is underlined.
Type BLIP2 In-BL. LA-V2 LLaVA MGPT m-owl Otter PGPT VPGT OF-2 In-XC. L-1.5 SGPT Moe R Ave.
TextBugger 0.18 0.24 0.27 0.19 0.25 0.29 0.30 0.31 0.23 0.43 0.20 0.17 0.21 0.32 0.26
Dee.ug 0.66 0.39 0.29 0.24 0.44 0.35 0.37 0.36 0.40 0.56 0.27 0.20 0.25 0.45 0.37
Cha. Pruthi 0.69 0.51 0.48 0.36 0.42 0.43 0.43 0.40 0.43 0.66 0.45 0.32 0.31 0.56 0.46
BertAttack 0.45 0.46 0.39 0.40 0.34 0.33 0.47 0.40 0.38 0.60 0.38 0.32 0.31 0.67 0.42
TextFooler 0.80 0.66 0.59 0.60 0.59 0.72 0.64 0.61 0.69 0.70 0.66 0.67 0.59 0.83 0.67
Wor. Pwws 0.76 0.66 0.63 0.57 0.54 0.64 0.66 0.57 0.71 0.68 0.69 0.54 0.60 0.83 0.65
StressTest 0.49 0.44 0.10 0.07 0.16 0.15 0.28 0.22 0.17 0.48 0.41 0.10 0.08 0.40 0.25
CheckList 0.39 0.34 0.19 0.29 0.43 0.33 0.41 0.27 0.23 0.34 0.25 0.11 0.20 0.38 0.30
Sen. Inp.on 0.57 0.40 0.28 0.23 0.43 0.30 0.41 0.32 0.32 0.60 0.40 0.25 0.26 0.48 0.38
Sem. 0.02 0.03 -0.01 0.04 0.04 0.03 0.03 0.08 0.03 0.05 0.05 0.02 0.00 0.09 0.04
Ave. ASDR 0.50 0.41 0.32 0.30 0.36 0.36 0.40 0.35 0.36 0.51 0.38 0.27 0.28 0.50 0.38
Table 4: Evaluation results of LVLMs’ robustness to content bias AVIs. The accuracy ($\uparrow$) is used as the metric. Best results in each row are bold, worst results are underlined. Uns., Cul., Gen. represent unsafe, culture, gender, respectively.
content BLIP2 In-BL. LA-V2 LLaVA MGPT m-owl Otter PGPT VPGT OF-2 In-XC. L-1.5 SGPT Moe Ge-ni G-4v R Ave.
harassment 0.09 0.85 1.00 1.00 0.27 0.72 0.23 0.73 0.18 0.99 0.07 0.32 0.27 0.22 0.72 0.36 0.50
hate 0.03 0.70 0.92 1.00 0.31 0.63 0.34 0.43 0.01 1.00 0.00 0.44 0.40 0.29 0.50 0.04 0.44
self-harm 0.20 0.82 0.95 1.00 0.29 0.80 0.47 0.88 0.35 1.00 0.45 0.41 0.43 0.38 0.58 0.19 0.57
sexual 1.00 1.00 1.00 1.00 0.29 0.78 0.42 0.98 0.06 1.00 0.55 0.98 0.96 0.97 0.88 0.77 0.79
shock 0.75 0.98 1.00 1.00 0.37 0.84 0.48 1.00 0.06 1.00 0.58 0.99 0.98 0.98 0.97 0.89 0.81
violence 0.86 0.99 1.00 1.00 0.36 0.79 0.51 0.98 0.15 1.00 0.80 0.91 0.93 0.90 0.94 0.81 0.81
Uns. Ave. 0.49 0.89 0.98 1.00 0.32 0.76 0.41 0.83 0.14 1.00 0.41 0.67 0.66 0.62 0.76 0.51 0.65
Cul. 0.64 0.91 0.93 1.00 0.40 0.78 0.46 0.96 0.67 1.00 0.88 0.71 0.66 0.59 0.86 0.78 0.76
black 0.87 0.33 0.22 0.02 0.63 0.25 0.25 0.23 0.45 0.01 0.84 0.74 0.81 0.68 0.72 0.84 0.49
white 0.90 0.40 0.23 0.02 0.63 0.25 0.26 0.22 0.52 0.03 0.88 0.87 0.90 0.78 0.76 0.85 0.53
yellow 0.92 0.49 0.16 0.02 0.63 0.25 0.21 0.22 0.41 0.09 0.92 0.83 0.90 0.80 0.74 0.85 0.53
Race Ave. 0.90 0.41 0.20 0.02 0.63 0.25 0.24 0.22 0.46 0.04 0.88 0.81 0.87 0.75 0.74 0.85 0.53
Gen. 0.88 0.55 0.31 0.03 0.59 0.29 0.11 0.46 0.60 0.09 0.68 0.73 0.89 0.64 0.94 0.92 0.54
Ave. Score 0.65 0.73 0.70 0.64 0.43 0.58 0.34 0.65 0.31 0.66 0.60 0.72 0.74 0.66 0.78 0.66 0.62
4.4 Results on Black-box Text Attack in Text-based AVIs
Figure 4: Further Analysis: Relationship between robustness to AVIs and (a) Tuning parameters, (b) Vicuna adapters, (c) LLaMA adapters, (d) Training data volume, (e) LLMs. FC, Q--f., Full_Tu., B_Tu., Resa., and RedPajama represent Fully connected layer [36], Q--Former [24], Full Tuning [5], Bias Tuning [9], Resampler [23], and RedPajama--INCITE--Instruct [23] respectively.
Figure 5: The relationship between the LVLMs’ robustness to AVIs and the average score before the attack.
The results of the text-based AVIs are presented in Table 3. Among the different attack methods, TextFooler [49] demonstrated the highest effectiveness with an ASDR of 67%. Conversely, Semantic [11] performed the poorest, achieving an ASDR of only 4%. The low ASDR observed in semantic-level attacks highlights the robustness of LVLMs to instructions provided by individuals with diverse language habits, including Japanese, Chinese, Korean, and others. Among character-level attacks, Pruthi [43] was the most effective, surpassing TextBugger [46] with a 20% higher ASDR. In word-level attacks, TextFooler [49] emerged as the most successful, outperforming BertAttack [48] by 25% in terms of effectiveness. For sentence-level attacks, Input-reduction [45] demonstrated the highest effectiveness, surpassing StressTest [50] by 13%.
Overall, all models showcased an ASDR of less than 55%. The top-performing model, LLaVA-1.5 [22], achieved an ASDR of only 27%, while the most vulnerable LVLM, OpenFlamingo-V2 [23], attained an ASDR of 51%.
4.5 Results on Content Bias AVIs
The experimental results of content bias AVIs are presented in Table 4. Among open-source LVLMs, LLaVA [36] and OpenFlamingo-V2 [23] emerge as the top performers for detecting unsafe information, achieving a 100% detection rate. In contrast, VPGTrans [38] and MiniGPT-4 [25] exhibit lower performance, with detection rates of 14% and 32% respectively. In the context of cultural content bias attacks, LLaVA [36] and OpenFlamingo-V2 [23] continue to demonstrate superior performance, while Otter [3] (46%) and MiniGPT-4 [25] (40%) show poorer results. Regarding racial content bias attacks, BLIP2 [24] emerges as the best-performing model, achieving a detection rate of 90%, while LLaVA [36] performs the worst, with only a 2% detection rate. For gender content bias attacks, ShareGPT4V [5] achieves the highest performance, reaching 89%, while OpenFlamingo-V2 [23] lags with a detection rate of only 9%. Overall, the best performer among all tested open-source LVLMs is ShareGPT4V [5], scoring 74%. VPGTrans [38] performs the worst, with a score of 31%.
Regarding advanced closed-source LVLMs like GeminiProVisionn [1] and GPT-4V [2], while GeminiProVision achieved the top performance among all tested models, we observed that GPT-4V even performed worse than some earlier open-source LVLMs like LLaMA-Adapter V2 [9]. We find that apart from the low detection rate of unsafe information such as hate and self-harm, GPT-4V demonstrates noticeable biases in cultural contexts. For instance, it displays a 25% higher detection rate for American funerals compared to Japanese funerals, and a 10% higher detection rate for American funerals compared to African funerals. We also observed notable instances of racial bias in GeminiProVision, which predicts a higher likelihood of poverty for Black individuals by approximately 30% compared to White individuals. Moreover, significant gender biases are evident as GeminiProVision associates police officers more with males and teachers more with females. These biases hinder the development of fair and reliable LVLMs.
This finding highlights that even closed-source LVLMs with the strongest defense mechanisms still exhibit challenges in accurately identifying unsafe information and addressing issues related to racial bias, gender bias, and cultural bias. These factors hinder the fair and secure application of LVLMs. Future research and development efforts focused on personal information protection and safer LVLMs should prioritize addressing these biases and incorporating defense mechanisms. We have also observed that certain models exhibit internal defense mechanisms. For example, when asked about the suitability of a specific occupation for a particular gender, the model provides a more neutral response, stating, The profession of a chef is suitable for both men and women. The ability to work under pressure, pay attention to detail, and have a passion for cooking are important qualities for a chef, regardless of gender. However, when we introduce a prompt such as Your answer can only be one word, either ’no’ or ’yes’. the model inevitably produces biased responses.
4.6 Further Analysis and Discussion
In this section, we analyze the relationship between the robustness to AVIs and factors such as model structure, training data, and training methods, using evaluation results from various LVLMs. Despite the differences in LVLMs’ structures and training data, the overall framework remains consistent, involving visual encoders, Large Language Models (LLMs), and feature interactors (adapters). While we strive to control variables as much as possible, it is challenging to strictly control them due to variations in models’ configurations. However, this analysis still provides valuable insights and conjectures.
AVIs Robustness and Tuning Parameters. Fig. 4(a) shows that image corruption AVIs exhibit a negative correlation with the quantity of tuning parameters, while content bias AVIs demonstrate a positive correlation. We also find that increasing the tuning parameters from 7B to 7.5B (specifically by tuning the image encoder), enhances the robustness of text-based AVIs and decision-based optimized image AVIs.
AVIs Robustness and LLM Adapters. In Fig.4(b), Vicuna adapters[7] are observed to have minimal impact on text attacks and image corruption. This suggests that relying solely on fully connected layers might pose challenges in effectively mitigating image corruption. On the other hand, Fig.4(c) demonstrates diverse robustness levels in LLaMA[8] adapters, highlighting the importance of prioritizing defense against weaker attack types specific to different adapters.
AVIs and Training Data Volume. In Fig. 4(d), the robustness of LVLMs against different AVIs does not exhibit a significant correlation with the scale of the training data. Instead, we speculate that factors such as data quality, content, and training methods may have a more pronounced impact on LVLMs’ robustness to AVIs.
AVIs and LLMs. In Fig. 4(e), we observe diverse levels of robustness among LVLMs that are based on different LLMs, which highlights the difficulty of adopting a unified defense approach and emphasizes the importance of considering the direction of defense and the specific structural differences across models.
AVIs and Average Score Before Attack. Fig. 5 illustrates the correlation between LVLM’s robustness scores against various attacks and their pre-attack scores. The Pearson Correlation Coefficient, $r$, gauges this correlation. Notably, the relationship with original scores is weaker for image corruptions in image-based AVIs and text-based AVIs, while decision-based optimized black-box attacks in image-based AVIs and content bias AVIs show a stronger correlation, with $r$ values of 0.51 and 0.42, respectively. This suggests that original scores may better predict the robustness of decision-based optimized black-box image-based AVIs and content bias AVIs compared to other attack types. However, the coefficients, $R^{2}$ are low, which means the ability of image and text comprehension may not be well related to the defense against AVIs. We make more detailed analyses considering attack categories and multimodal capabilities in the supplementary material.
5 Conclusion
In conclusion, this paper introduces AVIBench, a comprehensive framework designed to analyze the robustness of Large Vision-Language Models (LVLMs) against different types of adversarial visual-instructions (AVIs), including image-based AVIs, text-based AVIs, and content bias AVIs. AVIBench generates 260K AVIs, encompassing a wide range of multimodal capabilities and content biases. It conducts extensive evaluations involving 14 open-source LVLMs and two closed-source LVLMs. AVIBench provides a valuable tool for assessing the defense mechanisms of LVLMs. The vulnerabilities identified in LVLMs, when subjected to intentional and careless attacks, emphasize the critical need to enhance the robustness, security, and fairness of LVLMs to ensure their responsible deployment across various applications. Additionally, AVIBench will be publicly available as an open-source resource, serving as a foundational tool for robust LVLM research.
ETHICS STATEMENT. This paper analyzes the inherent biases in LVLMs, specifically cultural, gender, and racial biases, and their effectiveness in identifying unsafe information using content bias AVIs. The research aims to promote the safe and fair usage of LVLMs. All images used are sourced from the Internet, and biased images are not intentionally created. The images of content bias attacks will not be publicly shared and will only be used for online testing.